OpenStudy (anonymous):

Easy math, please help! Laura has a coin purse with eight coins in it. If she empties the purse onto the ground, what is the probability that all eight coins will show tails? A. 1/2 B. 1/4 C. 1/256 D. 1/128

OpenStudy (anonymous):

(1/2)(1/2)...(1/2) = (1/2)^8

OpenStudy (icalibear):

so there is a total of 8 coins, so 8 goes on the bottom of the fraction. Theoretically, there is a 50/50 chance that she will get tails, so 1/2 of 8 is 4 4/8 = 1/2

OpenStudy (anonymous):

All eight land tails is very unlikely!

OpenStudy (icalibear):

If this is theoretical probablility then its just a 50/50 chance

OpenStudy (anonymous):

No. Any one coin has 50/50 chance. To have all 8 land as tails is the product of the individual probabilities, thus (1/2)^8 = 1/256

OpenStudy (anonymous):

uhh... @icalibear @douglaswinslowcooper

OpenStudy (anonymous):

Imagine doing it. You shake the purse and the first one lands tails, then the second, then the third, then the fourth...wouldn't that seem very unusual?

OpenStudy (anonymous):

Yes, but you never know it could happen :-; @douglaswinslowcooper

OpenStudy (anonymous):

Yes, and you would not expect it half the times!

OpenStudy (anonymous):

True! @douglaswinslowcooper